OpenStudy (anonymous):

In a population, 93% of the people are right-handed, and 7% of the people are left-handed. In each of the last 100 days, 1 person in the population has been randomly stopped on the street, For the first 93 days, the person stopped was right-handed, and for the last 7 days, the person stopped was left-handed. If the same person is allowed to be stopped on more than 1 day, what is the probability that the next person stopped will be right-handed?

OpenStudy (anonymous):

@Yttrium

OpenStudy (btaylor):

There is an equal probability each day, since it is an independent event. Like if I were to flip a coin, knowing that my last 2, 20, or 2000 flips were heads does not change the likelihood that my next flip will be heads.

OpenStudy (anonymous):

@BTaylor so the probability is still 93%?

OpenStudy (btaylor):

Yes. Good job!

OpenStudy (anonymous):

Morgan is playing a board game that requires three standard dice to be thrown at one time. Each die has six sides, with one of the numbers 1 through 6 on each side. She has one throw of the dice left, and she needs a 17 to win the game. What is the probability that Morgan wins the game (order matters)? @BTaylor

OpenStudy (anonymous):

The only combination of numbers to make 17 is 5+6+6 right ?

OpenStudy (btaylor):

You can get a sum of 17 with a 6, a 6, and a 5. With different arrangements, there are 3 ways you could throw a 17. (5,6,6 6,5,6 6,6,5) There are \(6^3 = 216\) possibilities for the last throw. Can you turn this into a probability?

OpenStudy (anonymous):

1/6 ?

OpenStudy (anonymous):

noo , that can't be right because thats not one of my choices.

OpenStudy (anonymous):

1/216 ?

OpenStudy (btaylor):

it would be 3/216 because that is the number of target ways (to achieve 17) divided by the number of possible ways (216). Simplifying, it would be 1/72. Is that an answer choice?

OpenStudy (anonymous):

yes .

OpenStudy (anonymous):

Mary Katherine has a bag of 3 red apples, 5 yellow apples and 4 green apples. Mary takes a red apple out of the bag and does not replace it. What is the probability that the next apple she takes out is yellow?

OpenStudy (anonymous):

is that 5/11 ?

OpenStudy (btaylor):

yep!

OpenStudy (anonymous):

A quality inspector at a valve manufacturer randomly selects one valve from each batch of fifty valves to inspect for noncompliance. The first batch of fifty valves has five non-compliant ones. The second batch of fifty valves has four non-compliant valves. Find the probability that the inspector selects a non-compliant valve both times.

OpenStudy (anonymous):

can you help with that one ?

OpenStudy (btaylor):

P(first one AND second one) = P(first one) x P(second one) What is the probability each time? Multiply those (since they're independent, you can do that) and that is the probability that both are.

OpenStudy (anonymous):

I thought it would be 20/100 or 2/10 or 1/5 but none of those are a choice.

OpenStudy (btaylor):

you have: \[\frac{5}{50} \times \frac{4}{50} = \frac{1}{10} \times \frac{2}{25} = \frac{2}{250} = \frac{1}{125}\] I think you added the denominator instead of multiplying.

OpenStudy (anonymous):

Awhh man , you beat me to it I was just about to say that I got 1/125 the second time I did it .

OpenStudy (btaylor):

haha but you got it. I think you get it now!

OpenStudy (anonymous):

If events A and B are DEPENDENT, then A) A and B must occur together. B) A and B cannot occur together. C) A's occurrence can affect the probability of B's occurrence. D) A's occurrence cannot affect the probability of B's occurrence. *** C is correct right ? or A ?

OpenStudy (btaylor):

C is correct.